OpenStudy (devonhoward15):

which equation represents a direct linear variation?... will fan and medal A.y=x^2 B.y=1/3x C.y=x-3 D.y=1/x

OpenStudy (photon336):

linear variation pretty sure that means a linear equation in the form y = m(x)+b

OpenStudy (devonhoward15):

yes

OpenStudy (photon336):

so which one would you think it is based on what I said? @Devonhoward15

OpenStudy (devonhoward15):

c

OpenStudy (devonhoward15):

@Photon336

OpenStudy (photon336):

yes

OpenStudy (devonhoward15):

thank you

OpenStudy (devonhoward15):

could you help me with one more?

OpenStudy (devonhoward15):

which equation is the direct linear variation equation for the relationship? A.y=x+9 B.y=x-9 C.y=4x D.y=1/4x

OpenStudy (devonhoward15):

i think its a

OpenStudy (devonhoward15):

@Photon336

OpenStudy (photon336):

that's linear

OpenStudy (devonhoward15):

yeah so is it not a?